# Proxy-GS: Efficient 3D Gaussian Splatting via Proxy Mesh

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as an efficient approach for achieving photorealistic rendering. Recent MLP-based variants further improve visual fidelity but introduce substantial decoding overhead during rendering. To alleviate computation cost, several pruning strategies and level-of-detail (LOD) techniques have been introduced, aiming to effectively reduce the number of Gaussian primitives in large-scale scenes. However, our analysis reveals that significant redundancy still remains due to the lack of occlusion awareness. In this work, we propose Proxy-GS, a novel pipeline that exploits a proxy to introduce Gaussian occlusion awareness from any view. At the core of our approach is a fast proxy system capable of producing precise occlusion depth maps at resolution $\times1000$ under $1\text{\,}\mathrm{m}\mathrm{s}$. This proxy serves two roles: first, it guides the culling of anchors and Gaussians to accelerate rendering speed. Second, it guides the densification towards surfaces during training, avoiding inconsistencies in occluded regions, and improving the rendering quality. In heavily occluded scenarios, such as the MatrixCity Streets dataset, Proxy-GS not only equips MLP-based Gaussian splatting with stronger rendering capability but also achieves faster rendering speed than the original 3DGS. Specifically, it achieves more than $2.5\times$ speedup over Octree-GS, and consistently delivers substantially higher rendering quality. Code will be public upon acceptance.
@@ 1 Introduction
With the emergence of Neural Radiance Fields (NeRF) (Mildenhall et al., 2020), high-quality novel view synthesis has become possible, but the slow rendering speed limits its practical use. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023b) has significantly improved efficiency, greatly advancing AR and VR applications (Xu et al., 2024; Zhan et al., 2024; Niu et al., 2025). However, vanilla 3DGS often produces heavily redundant Gaussians that attempt to fit every training view while neglecting the underlying scene geometry. To address this limitation and pursue higher-fidelity representations, structured MLP-based Gaussian approaches such as scaffold-GS (Lu et al., 2024) and Octree-GS (Ren et al., 2024) have been introduced.
At the core of the MLP-based 3DGS method lies an MLP decoder that conditions on the camera viewing direction to dynamically generate Gaussian attributes. Although these structured Gaussian variants substantially strengthen the modeling of challenging and view-dependent details, they also impose extra decoding operations at inference, leading to increased computational cost. This drawback becomes particularly critical in large-scale scene reconstruction (Liu et al., 2024; Lin et al., 2024; Gao et al., 2025; Li et al., 2024a; Gao et al., 2024), where the number of Gaussian primitives and rendering complexity grow dramatically, making efficient decoding and rendering indispensable.
Although pruning strategies (Fan et al., 2024; Lee et al., 2024; Liu et al., 2025) can be introduced to reduce redundancy, they inevitably lead to a loss in rendering quality. Meanwhile, following works (Ren et al., 2024; Kerbl et al., 2024; Cui et al., 2024) employ a level-of-detail (LOD) structure to mitigate redundancies from distant scene contents, but this approach is mainly effective in relatively occlusion-free environments. In contrast, real-world scenarios are full of occlusions, especially in large-scale modern city streets and complex indoor environments with multiple rooms. For future ultra-large VR walkthroughs that seamlessly span from indoor to outdoor scenes, effective occlusion culling becomes an essential and intuitive requirement.
Moreover, since most practitioners rely on consumer-grade GPUs rather than datacenter-oriented ones such as A100s, it is important to consider the hardware characteristics of these devices. Consumer GPUs, typically designed for gaming and graphics applications, are equipped with dedicated hardware rasterization units. The widespread adoption of 3DGS thus requires careful adaptation to hardware rasterization (fas, 2024).
@@
As shown in Fig. 1, Proxy-GS not only achieves up to a $3\times$ speedup in rendering on top of existing MLP-based LOD frameworks Octree-GS (Ren et al., 2024) but also improves occlusion awareness in anchor selection, leading to higher rendering quality. Our main contributions can be summarized as follows:
@@
  • We leverage engineering optimizations to reduce the time of acquiring a $1000^{2}$-resolution depth map to under $1\text{\,}\mathrm{m}\mathrm{s}$.
@@ 2.1 Neural Rendering
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) pioneered the idea of representing a scene as a volumetric radiance field, enabling high-quality novel view synthesis for bounded scenes, typically centered around a single object. Subsequent extensions improved the scalability and visual fidelity of NeRF-based methods: Mip-NeRF (Barron et al., 2021) introduced proper anti-aliasing to handle multi-scale observations, NeRF++ (Zhang et al., 2020) lifted the constraint of strictly bounded scenes, and Mip-NeRF 360 (Barron et al., 2022) extended anti-aliased representations to unbounded, object-centric settings. Despite these advances, NeRF-style volumetric rendering remains computationally expensive due to the need for dense ray sampling and neural field evaluation.
@@
To overcome this inefficiency, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023a) was recently proposed as an explicit point-based alternative. By representing a scene with a set of anisotropic 3D Gaussian primitives and employing a splatting-based rasterization pipeline, 3DGS enables real-time rendering while preserving high visual quality. This paradigm shift bridges the gap between neural radiance fields and traditional graphics pipelines, offering both efficiency and scalability. While 3DGS achieves real-time performance with explicit Gaussian primitives, its reliance on directly optimized parameters often leads to limited expressiveness, particularly in capturing fine-grained appearance details and complex view-dependent effects. To address this shortcoming, MLP-based extensions such as Scaffold-GS (Lu et al., 2024) and Octree-GS (Ren et al., 2024) introduce neural decoders that generate Gaussian attributes from learned anchor features. By leveraging structured anchors and neural decoding, these approaches significantly improve the representational capacity, enabling more accurate modeling of geometry and appearance in large and challenging scenes.
However, this enhanced expressiveness comes at the cost of efficiency. The dependence on per-anchor MLP decoding introduces substantial computational overhead during inference, making rendering speed a critical bottleneck. Even with level-of-detail (LOD) designs to reduce the number of anchors processed per view, MLP-based methods still struggle to balance quality and efficiency, especially in large-scale urban or indoor environments with heavy occlusions. To this end, we are the first to address such occlusion-induced redundancy through a lightweight proxy mechanism.
2.2 Toward Faster 3D Gaussian Splatting Rendering
@@
For rendering acceleration, many studies (Lee et al., 2024; Fan et al., 2024; Liu et al., 2025; Wang & Xu, 2025) have explored pruning or compression strategies to reduce the number of Gaussians and thus alleviate computational overhead. While such pruning-based methods can be effective to some extent, they inevitably face scalability bottlenecks in large scenes, where aggressive pruning results in performance degradation. Beyond pruning strategy, another line of research focuses on architectural designs for rendering acceleration. Among them, level-of-detail (LOD) architectures have become particularly influential. Hierarchical-GS (Kerbl et al., 2024) merges neighboring Gaussians to reduce rendering cost, achieving higher frame rates at the expense of some visual fidelity. LetsGo (Cui et al., 2024) jointly optimizes multi-resolution Gaussian models and demonstrates strong performance in LiDAR-based scenarios, yet its reliance on multi-resolution point cloud inputs incurs substantial training overhead and creates a strong dependence on point cloud accuracy. CityGaussian (Liu et al., 2024) further combines pruning strategies (Fan et al., 2024) with LOD-based rendering to enhance scalability in urban scenes.
@@
While the aforementioned works improve efficiency for explicit 3DGS, LOD mechanisms have also been extended to MLP-based Gaussians. Octree-GS (Ren et al., 2024) organizes anchors into a multi-level octree, where the level selection is determined by the distance to the camera, thereby reducing the number of anchors decoded at each frame. This strategy alleviates part of the computational burden in large-scale scenes, but the rendering speed still leaves considerable room for improvement. Recent work Cache-GS (Tao et al., 2025) provides further acceleration by reusing decoded Gaussians, effectively doubling the rendering speed of Octree-GS, although this comes with a noticeable loss in rendering quality. In parallel, methods like FLASH-GS (Feng et al., 2024) target low-level CUDA optimizations of the original 3DGS pipeline, aiming to improve efficiency at the kernel level. Recent work has also explored leveraging occlusion for accelerating rendering. For example,  Ye et al. (2025) proposed using pre-rendered depth maps to guide 3DGS rendering. However, their depth acquisition relies on 2DGS rendering, which is less efficient compared to our lightweight proxy-based approach.
@@ 3.1 MLP-based 3DGS
To exploit the structural priors provided by Structure-from-Motion (SfM), a line of work such as Scaffold-GS (Lu et al., 2024) and Octree-GS (Ren et al., 2024) has been developed. Instead of reconstructing Gaussians directly from sparse SfM points, Scaffold-GS first builds a coarse voxel grid and places anchor points at the voxel centers. Each anchor is associated with a latent feature vector $f$, which is fed into a multi-layer perceptron (MLP) to decode the corresponding Gaussian attributes:
@@
where $\theta$ denotes the MLP parameters, and $\mu_{j}$, $\Sigma_{j}$, $c_{j}$, and $\alpha_{j}$ represent the mean, covariance, color, and opacity of the $j$-th Gaussian derived from the $i$-th anchor under viewing direction $v_{i}$. The generated neural Gaussians are subsequently rasterized in the same way as explicit 3D Gaussians.
The advantage of anchor-based placement is that the decoded Gaussians inherit structural cues from the underlying SfM prior, which reduces redundancy and improves robustness for novel view rendering. Octree-GS extends this framework by substituting the voxel grid with an explicit octree representation, enabling the scene to be modeled at multiple resolutions.
The hierarchical design of the octree naturally supports level-of-detail (LOD) construction. During rendering, appropriate LOD levels can be selected adaptively based on the camera distance, thereby reducing decoding cost and improving scalability to larger-scale scenes.
@@ 4.1 Motivation
Existing MLP-based works, such as Octree-GS Ren et al. (2024) and Scaffold-GS Lu et al. (2024), design anchor structures to better exploit the inherent hierarchy and structural priors. However, since their anchor selection does not explicitly account for occlusions, the anchors are optimized merely to fit RGB images. As a result, the binding between anchors and their associated Gaussians can become inconsistent in space, leading to redundant decoding and degraded structural interpretability
@@ 4.2 Proxy Guided Filter
A central question in our study is how to obtain occlusion relationships both efficiently and with negligible loss of accuracy. We find that leveraging lightweight proxy meshes for hardware rasterization enables depth rendering at only a marginal time cost. For many outdoor large-scale scenes, dense point clouds are already available or can be generated using tools such as COLMAP. In contrast, indoor scenes often contain texture-less regions that cause SfM-based reconstruction to fail, so we adopt a hybrid strategy. Specifically, we combine state-of-the-art monocular depth estimation Wang et al. (2025) with PGSR Chen et al. (2024), in a manner similar to recent indoor surface reconstruction approaches Zhang et al. (2024); Ren et al. (2025). Further implementation details are provided in the Appendix A.4. We construct proxy meshes using existing engineering pipelines and apply surface simplification to retain only coarse geometric structures. This proxy is sufficient to fully exploit the high throughput of hardware fixed-function units for efficient depth generation.
To further accelerate the process, the scene is partitioned into fine-grained clusters, and hierarchical visibility checks such as Hierarchical Z-buffer (Hi-Z) culling Greene et al. (1993) are employed to quickly cull invisible clusters. In the fragment stage, early-fragment tests (Early-Z) are enabled, and we keep the fragment shader minimal by removing operations unrelated to depth writes. This allows our method to output depth maps at a high speed even in complex and large-scale urban scenes, as shown in Fig. 3. The depth map is kept on GPU and directly exploited in CUDA occlusion culling to avoid GPU-CPU-GPU round-trip overhead. For more details, please refer to the Appendix A.5.
Then we fuse the occlusion culling and frustum culling of anchors in a single CUDA kernel: Given an original anchor point $\mathbf{p}_{\text{orig}}=(x,y,z)\in\mathbb{R}^{3}$, the point is first transformed into the camera (view) coordinate system via the view matrix $V\in\mathbb{R}^{4\times 4}$:
$\mathbf{p}_{\text{view}}=V\begin{bmatrix}\mathbf{p}_{\text{orig}}\\ 1\end{bmatrix}$. (2)
Then it transfers to the homogeneous clip space using the projection matrix $P\in\mathbb{R}^{4\times 4}$:
$\mathbf{p}_{\text{hom}}=P\mathbf{p}_{\text{view}}=(x_{h},y_{h},z_{h},w_{h})^{\top}.$ (3)
To obtain normalized device coordinates (NDC), we divide by the homogeneous component:
$\mathbf{p}_{\text{ndc}}=\left(\frac{x_{h}}{w_{h}+\epsilon},\frac{y_{h}}{w_{h}+\epsilon},\frac{z_{h}}{w_{h}+\epsilon}\right),\quad\epsilon=10^{-7}.$ (4)
We denote the resulting coordinates as $(x_{\text{ndc}},y_{\text{ndc}},z_{\text{ndc}})$.
@@
Figure 3: Comparison of the time proportion of each inference component (Rendering, anchor filter, depth rendering) with that of Octree-GS.
@@ 4.3 Proxy-Guided Densification
In the original anchor-growing densification strategy, new anchors are generated around Gaussian splats that exhibit large gradients during training. However, this procedure may introduce redundant anchors behind the proxy mesh depth: although these Gaussians have large gradients, the newly grown anchors do not contribute to rendering due to occlusion.
To tackle this limitation, and inspired by the multi-view depth densification strategy in Li et al. (2024b), we introduce proxy-guided densification, which explicitly projects anchors onto the surface of the proxy mesh. Since proxy depth maps are pre-computed, we can measure the patch-wise L1 loss and identify regions where the rendering error is consistently large.
@@
For each selected patch $\mathcal{P}$, choose a representative pixel $(u_{\mathcal{P}},v_{\mathcal{P}})$ (e.g., the patch center), read the hardware depth $z_{h}(u_{\mathcal{P}},v_{\mathcal{P}})$, and convert it to linear camera-space depth with near/far $(n,f)$, to obtain $d_{\text{mesh}}(u_{\mathcal{P}},v_{\mathcal{P}})$. We then back-project this pixel to 3D and take it as the new anchor position:
@@
where $\kappa[\cdot]\in\mathbb{N}$ tracks the current number of anchors in each cell.
Table 1: Quantitative results on MatrixCity (Li et al., 2023). We report average results over Block 1&2, Block 3&4, and Block 5. (Block 1&2 and 3&4 represent the average evaluation metrics of their respective two blocks.) The best and second-best are highlighted.
Block 1&2 Block 3&4 Block 5
[…]
3DGS (Kerbl et al., 2023a) 21.55 0.730 0.366 115 20.78 0.739 0.372 114 20.70 0.697 0.425 121
Scaffold-GS (Lu et al., 2024) 21.44 0.721 0.375 81 20.56 0.727 0.376 66 20.56 0.693 0.426 71
Hierarchical-GS (Kerbl et al., 2024) 20.50 0.707 0.418 61 20.38 0.719 0.422 41 20.22 0.673 0.463 60
Octree-GS (Ren et al., 2024) 21.94 0.737 0.347 32 20.95 0.743 0.354 30 21.41 0.731 0.375 48
Table 2: Quantitative results on real world Outdoor and Indoor datasets (Xiong et al., 2024; Kerbl et al., 2024; Barron et al., 2023). best and second-best are highlighted.
CUHK-LOWER Berlin Small City
[…]
3DGS (Kerbl et al., 2023a) 25.48 0.729 0.389 138 27.79 0.907 0.223 187 22.90 0.727 0.372 132
Scaffold-GS (Lu et al., 2024) 26.30 0.785 0.282 117 27.80 0.912 0.213 128 20.00 0.713 0.370 62
Hierarchical-GS (Kerbl et al., 2024) 25.18 0.707 0.408 90 27.65 0.902 0.228 145 22.07 0.728 0.377 89
Hierarchical-GS($\tau_{1}$) 25.19 0.708 0.408 82 27.65 0.901 0.229 150 22.07 0.728 0.377 90
Hierarchical-GS($\tau_{2}$) 25.14 0.705 0.411 96 27.60 0.899 0.232 152 22.07 0.728 0.378 106
Hierarchical-GS($\tau_{3}$) 24.58 0.678 0.435 120 27.34 0.890 0.244 160 22.02 0.722 0.386 119
Octree-GS (Ren et al., 2024) 26.42 0.794 0.267 212 27.83 0.911 0.218 263 23.03 0.731 0.355 51

### v3

# Proxy-GS: Unified Occlusion Priors for Training and Inference in Structured 3D Gaussian Splatting

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as an efficient approach for achieving photorealistic rendering. Recent MLP-based variants further improve visual fidelity but introduce substantial decoding overhead during rendering. To alleviate computation cost, several pruning strategies and level-of-detail (LOD) techniques have been introduced, aiming to effectively reduce the number of Gaussian primitives in large-scale scenes. However, our analysis reveals that significant redundancy still remains due to the lack of occlusion awareness. In this work, we propose Proxy-GS, a novel pipeline that exploits a proxy to introduce Gaussian occlusion awareness from any view. At the core of our approach is a fast proxy system capable of producing precise occlusion depth maps at resolution $\times1000$ under $1\text{\,}\mathrm{m}\mathrm{s}$. This proxy serves two roles: first, it guides the culling of anchors and Gaussians to accelerate rendering speed. Second, it guides the densification towards surfaces during training, avoiding inconsistencies in occluded regions, thus improving the rendering quality. In heavily occluded scenarios such as the MatrixCity Streets dataset, Proxy-GS achieves more than $2.5\times$ speedup over Octree-GS while also improving rendering quality.
@@ 1 Introduction
With the emergence of Neural Radiance Fields (NeRF) [28], high-quality novel view synthesis has become possible, but the slow rendering speed limits its practical use. Recently, 3D Gaussian Splatting (3DGS) [15] has significantly improved efficiency, greatly advancing AR and VR applications [35, 38, 29]. However, vanilla 3DGS often produces heavily redundant Gaussians that attempt to fit every training view while neglecting the underlying scene geometry. To address this limitation and pursue higher-fidelity representations, structured MLP-based Gaussian approaches such as scaffold-GS [27] and Octree-GS [30] have been introduced.
At the core of the MLP-based 3DGS method lies an MLP decoder that conditions on the camera viewing direction to dynamically generate Gaussian attributes. Although these structured Gaussian variants substantially strengthen the modeling of challenging and view-dependent details, they also impose extra decoding operations at inference, leading to increased computational cost. This drawback becomes particularly critical in large-scale scene reconstruction [25, 23, 9, 20, 8], where the number of Gaussian primitives and rendering complexity grow dramatically, making efficient decoding and rendering indispensable.
Although pruning strategies [5, 19, 26] can be introduced to reduce redundancy, they inevitably lead to a loss in rendering quality. Meanwhile, following works [30, 16, 4] employ a level-of-detail (LOD) structure to mitigate redundancies from distant scene contents, but this approach is mainly effective in relatively occlusion-free environments. In contrast, real-world scenarios are full of occlusions, especially in large-scale modern city streets and complex indoor environments with multiple rooms. For future ultra-large VR walkthroughs that seamlessly span from indoor to outdoor scenes, effective occlusion culling becomes an essential and intuitive requirement.
Moreover, since most practitioners rely on consumer-grade GPUs rather than datacenter-oriented ones such as A100s, it is important to consider the hardware characteristics of these devices. Consumer GPUs, typically designed for gaming and graphics applications, are equipped with dedicated hardware rasterization units. The widespread adoption of 3DGS thus requires careful adaptation to hardware rasterization [6].
@@
As shown in Fig. 1, Proxy-GS not only achieves up to a $3\times$ speedup in rendering on top of existing MLP-based LOD frameworks Octree-GS [30] but also improves occlusion awareness in anchor selection, leading to higher rendering quality. Our main contributions can be summarized as follows:
@@
  • We leverage hardware rasterization to reduce the time of acquiring a $1000^{2}$-resolution depth map to under $1\text{\,}\mathrm{m}\mathrm{s}$.
@@ 2.1 Neural Rendering
Neural Radiance Fields (NeRFs) [28] pioneered the idea of representing a scene as a volumetric radiance field, enabling high-quality novel view synthesis for bounded scenes, typically centered around a single object. Subsequent extensions improved the scalability and visual fidelity of NeRF-based methods: Mip-NeRF [1] introduced proper anti-aliasing to handle multi-scale observations, NeRF++ [39] lifted the constraint of strictly bounded scenes, and Mip-NeRF 360 [3] extended anti-aliased representations to unbounded, object-centric settings. Despite these advances, NeRF-style volumetric rendering remains computationally expensive due to the need for dense ray sampling and neural field evaluation.
@@
To overcome this inefficiency, 3D Gaussian Splatting (3DGS) [14] was recently proposed as an explicit point-based alternative. While 3DGS achieves real-time performance with explicit Gaussian primitives, its reliance on directly optimized parameters often leads to limited expressiveness, particularly in capturing fine-grained appearance details and complex view-dependent effects. To address this shortcoming, MLP-based extensions such as Scaffold-GS [27] and Octree-GS [30] introduce neural decoders that generate Gaussian attributes from learned anchor features. By leveraging structured anchors and neural decoding, these approaches significantly improve the representational capacity, enabling more accurate modeling of geometry and appearance in large and challenging scenes.
2.2 Efficient 3D Gaussian Splatting Rendering
@@
For rendering acceleration, many studies [19, 5, 26, 33] have explored pruning or compression strategies to reduce the number of Gaussians and thus alleviate computational overhead. While such pruning-based methods can be effective to some extent, they inevitably face scalability bottlenecks in large scenes, where aggressive pruning results in performance degradation. Beyond the pruning-based strategy, another line of research focuses on architectural designs for rendering acceleration. Among them, level-of-detail (LOD) architectures have become particularly influential. Hierarchical-GS [16] merges neighboring Gaussians to reduce rendering cost, achieving higher frame rates at the expense of some visual fidelity. LetsGo [4] jointly optimizes multi-resolution Gaussian models and demonstrates strong performance in LiDAR-based scenarios, yet its reliance on multi-resolution point cloud inputs incurs substantial training overhead and creates a strong dependence on point cloud accuracy. CityGaussian [25] further combines pruning strategies [5] with LOD-based rendering to enhance scalability in urban scenes. LODGE [17] proposes an LOD framework that adapts visible Gaussians across scales for memory and speed-constrained rendering. Horizon-GS [12] unifies aerial-to-ground reconstruction and rendering with scalable data organization for large environments. Virtualized 3D Gaussians [36] introduces a cluster-based LOD with hierarchical Gaussian groups and online selection for composed scenes, and VastGaussian [23] demonstrates city-scale reconstruction via progressive partition and merging.
@@
While the aforementioned works improve efficiency for explicit 3DGS, more LOD mechanisms have also been extended to MLP-based Gaussians. Octree-GS [30] organizes anchors into a multi-level octree, where the level selection is determined by the distance to the camera, thereby reducing the number of anchors decoded at each frame. This strategy alleviates part of the computational burden in large-scale scenes, but the rendering speed still leaves considerable room for improvement. Recent work Cache-GS [31] provides acceleration by reusing decoded Gaussians, doubling the rendering speed of Octree-GS, although this comes with a noticeable loss in rendering quality. In parallel, methods like FLASH-GS [7] target low-level CUDA optimizations of the original 3DGS pipeline, aiming to improve efficiency at the kernel level. While OccluGaussian [24] reasons about occlusion by partitioning scenes into clusters, our approach performs per-pixel, proxy-guided filtering, which preserves fine details and aligns with actual rendering cost. Recent work has also explored leveraging occlusion for accelerating rendering. For example,  Ye et al. [37] proposed using pre-rendered depth maps to guide 3DGS rendering. However, their depth acquisition relies on surfel rendering, which is less efficient compared to our lightweight proxy-based approach.
@@ 3.1 MLP-based 3DGS
To exploit the structural priors provided by Structure-from-Motion (SfM), a line of work such as Scaffold-GS [27] and Octree-GS [30] has been developed. Instead of reconstructing Gaussians directly from sparse SfM points, Scaffold-GS first builds a coarse voxel grid and places anchor points at the voxel centers. Each anchor is associated with a latent feature vector $f$, which is fed into a multi-layer perceptron (MLP) to decode the corresponding Gaussian attributes:
@@
where $\theta$ denotes the MLP parameters, and $\mu_{j}$, $\Sigma_{j}$, $c_{j}$, and $\alpha_{j}$ represent the mean, covariance, color, and opacity of the $j$-th Gaussian derived from the $i$-th anchor under viewing direction $v_{i}$. The generated neural Gaussians are subsequently rasterized in the same way as explicit 3D Gaussians. The advantage of anchor-based placement is that the decoded Gaussians inherit structural cues from the underlying SfM prior, which reduces redundancy and improves robustness for novel view rendering. Octree-GS extends this framework by substituting the voxel grid with an explicit octree representation, enabling the scene to be modeled at multiple resolutions. The hierarchical design of the octree naturally supports level-of-detail (LOD) construction. During rendering, appropriate LOD levels can be selected adaptively based on the camera distance, thereby reducing decoding cost and improving scalability to larger-scale scenes.
@@ 4.1 Motivation
Existing MLP-based works, such as Octree-GS [30] and Scaffold-GS [27], design anchor structures to better exploit the inherent hierarchy and structural priors. However, since their anchor selection does not explicitly account for occlusions, the anchors are optimized merely to fit RGB images. As a result, the binding between anchors and their associated Gaussians can become inconsistent in space, leading to redundant decoding and degraded structural interpretability.
@@ 4.2 Proxy Guided Filter
A central question in our study is how to obtain occlusion relationships both efficiently and with negligible loss of accuracy. We find that leveraging lightweight proxy meshes for hardware rasterization enables depth rendering at only a marginal time cost. For many outdoor large-scale scenes, dense point clouds are already available or can be generated using tools such as COLMAP. In contrast, indoor scenes often contain texture-less regions that cause SfM-based reconstruction to fail. As large reconstruction models [32] already achieve very strong performance, especially for indoor scenes. In our pipeline, we leverage MapAnything [13], using Colmap pose and RGB image as input to obtain a dense point cloud for indoor environments and then convert it into a mesh. Further implementation details are provided in the Appendix 7.5. We construct proxy meshes using existing different pipelines and apply surface simplification to retain only coarse geometric structures. This proxy is sufficient to fully exploit the high throughput of hardware fixed-function units for efficient depth generation.
As shown in Fig.  2, to further accelerate the process, the mesh is partitioned into fine-grained clusters, and hierarchical visibility checks such as Hierarchical Z-buffer (Hi-Z) culling [10] are employed to quickly cull invisible clusters. In the fragment stage, Early-Z is enabled, and we keep the fragment shader minimal by removing operations unrelated to depth writes. This allows our method to output depth maps at a high speed even in complex and large-scale urban scenes, as shown in Fig. 3. The depth map is kept on GPU and directly exploited in CUDA occlusion culling to avoid GPU-CPU-GPU round-trip overhead. We also compare the speed of our method with other conventional depth-acquisition approaches in the Appendix  LABEL:depth-acquisition, further demonstrating our efficiency advantages.
Then we fuse the occlusion culling and frustum culling of anchors in a single CUDA kernel: We denote the pixels ndc coordinates as $(x_{\text{ndc}},y_{\text{ndc}},z_{\text{ndc}})$.
@@
Figure 3: Comparison of the time proportion of each inference component (Rendering, anchor filter, depth rendering) with that of Octree-GS on MatrixCity dataset.
@@ 4.3 Proxy-Guided Densification
In the original anchor-growing densification strategy, new anchors are generated around Gaussian splats that exhibit large gradients during training. However, this procedure may introduce redundant anchors behind the proxy mesh depth: although these Gaussians have large gradients, the new anchors feature do not decode due to occlusion.
To tackle this limitation, and inspired by the multi-view depth densification strategy in [22], we introduce proxy-guided densification, which explicitly projects anchors onto the surface of the proxy mesh. Since proxy depth maps are pre-computed, we can measure the patch-wise L1 loss and identify regions where the rendering error is consistently large.
Table 1: Quantitative results on MatrixCity [21]. We report average results over Block 1&2, Block 3&4, and Block 5. (Block 1&2 and 3&4 represent the average evaluation metrics of their respective two blocks.) The best and second-best are highlighted.
Block 1&2 Block 3&4 Block 5
[…]
3DGS [14] 21.55 0.730 0.366 115 20.78 0.739 0.372 114 20.70 0.697 0.425 121
Scaffold-GS [27] 21.44 0.721 0.375 81 20.56 0.727 0.376 66 20.56 0.693 0.426 71
Hierarchical-GS [16] 20.50 0.707 0.418 61 20.38 0.719 0.422 41 20.22 0.673 0.463 60
Octree-GS [30] 21.94 0.737 0.347 32 20.95 0.743 0.354 30 21.41 0.731 0.375 48
Table 2: Quantitative results on real world outdoor and indoor datasets [34, 16, 2]. The best and second-best are highlighted. Small City has more severe occlusions, while Berlin and CUHK-LOWER have relative weaker occlusions.
Small City Berlin CUHK-LOWER
[…]
3DGS [14] 22.90 0.727 0.372 132 27.79 0.907 0.223 187 25.48 0.729 0.389 138
Scaffold-GS [27] 20.00 0.713 0.370 62 27.80 0.912 0.213 128 26.30 0.785 0.282 117
Hierarchical-GS [16] 22.07 0.728 0.377 89 27.65 0.902 0.228 145 25.18 0.707 0.408 90
Hierarchical-GS($\tau_{1}$) 22.07 0.728 0.377 90 27.65 0.901 0.229 150 25.19 0.708 0.408 82
Hierarchical-GS($\tau_{2}$) 22.07 0.728 0.378 106 27.60 0.899 0.232 152 25.14 0.705 0.411 96
Hierarchical-GS($\tau_{3}$) 22.02 0.722 0.386 119 27.34 0.890 0.244 160 24.58 0.678 0.435 120
Octree-GS [30] 23.03 0.731 0.355 51 27.83 0.911 0.218 263 26.42 0.794 0.267 212
Proxy-GS 23.09 0.736 0.344 139 27.85 0.912 0.216 275 26.44 0.795 0.262 239
@@
For each selected patch $\mathcal{P}$, we choose a pixel $(u_{\mathcal{P}},v_{\mathcal{P}})$ (e.g., the patch center), read the hardware depth $z_{h}(u_{\mathcal{P}},v_{\mathcal{P}})$, and convert it to linear camera-space depth with near/far $(n,f)$, to obtain $d_{\text{mesh}}(u_{\mathcal{P}},v_{\mathcal{P}})$. We then back-project this pixel to 3D and take it as the new anchor position:
@@
where $\kappa[\cdot]\in\mathbb{N}$ counts the anchors in each cell.